# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ 2.1 Risk estimation
Quantitative risk measurement often relies on precise estimation of risk measures commonly used in finance and actuarial science (McNeil et al.,, 2005). Calculating risk for multidimensional random variables can be challenging due to the need for complex integrals, often approximated using Monte Carlo simulation. When the underlying distribution isn’t directly accessible and only limited samples are available, Monte Carlo-based risk estimators tend to underestimate the actual risk. Kim and Hardy, (2007) addressed this by using bootstrapping to correct bias in Value at Risk (VaR) and Conditional Tail Expectation (CTE) estimates, with Kim, (2010) extending this approach to general distortion risk measures. Bootstrapping involves sampling with replacement from the empirical distribution, calculating the statistic, and averaging the outcomes over multiple iterations. However, bootstrap estimates still tend to underestimate risk in finite samples due to the lack of extreme tail scenarios (see Figure 4). In contrast, our bootstrapping procedure draws samples from a ”bias-aware” fitted distribution that better accounts for tail scenarios.
Several approaches have been proposed in the literature for estimating tail risks. Lam and Mottet, (2017) introduce a distributionally robust optimization based method to construct worst-case bounds on tail risk, assuming the density function is convex beyond a certain threshold. Extreme Value Theory (EVT) is commonly used to estimate tail risk measures, such as CVaR, by fitting a Generalized Pareto Distribution to values exceeding a threshold. Troop et al., (2021) develop an asymptotically unbiased CVaR estimator by correcting the bias in the estimates obtained via maximum likelihood. While these methods focus on upper-tail risk, they are not directly applicable to the entropic risk measure. Instead of fitting an extreme value distribution, we utilize a parametric two-component Gaussian Mixture Model (GMM), which provides a closed-form expression for the entropic risk and captures both the mean and the tails of the data.
@@ 2.2 Correcting optimistic bias
Several approaches have been proposed to correct the bias of SAA in linear optimization problems, assuming the true data distribution is Gaussian. Ito et al., (2018) propose a perturbation method to obtain an asymptotically unbiased estimator of the true loss, generating parameters around the true one under the assumption of Gaussian error. Similarly, Gupta et al., (2022) provide a procedure to debias the in-sample performance of affine policies under a Gaussian assumption. However, extending these methods to our non-linear problem is challenging, and our model isn’t limited to Gaussian distributions.
In Siegel and Wagner, (2023), the authors analytically characterize the bias in SAA policies for a data-driven newsvendor problem, providing an asymptotically debiased profit estimator by leveraging the asymptotic properties of order statistics. Iyengar et al., (2023) introduce an Optimizer’s Information Criterion (OIC) to correct bias in SAA policies, generalizing the approach by Siegel and Wagner, (2023). However, OIC requires access to the gradient, Hessian, and influence function of the decision rule, which can be challenging to obtain in general constrained optimization problems. Moreover, the form of optimal policy is known for risk-neutral newsvendor problems but not for entropic risk minimization problems.
@@ 2.3 Insurance pricing
The design of insurance contracts has been widely studied since the foundational work of Arrow (Arrow,, 1963, 1971). Under the assumption that premiums are proportional to the policy’s actuarial value, it has been shown that an expected utility-maximizing policyholder will choose full coverage above a deductible. Various extensions of Arrow’s model have been proposed to account for the risk aversion of both the insured and the insurer, using criteria such as mean-variance (Kaluszka, 2004a, ; Kaluszka, 2004b, ), Value at Risk (VaR), and Tail VaR (Cai et al.,, 2008). Bernard and Tian, (2010) incorporate regulatory constraints on the insurer’s insolvency risk through VaR. Cheung et al., (2014) extend these models to multiple policyholders with fully dependent risks (comonotonicity), where the insurer utilizes convex law-invariant risk measures. Bernard et al., (2020) further explore different levels of dependence among policyholders, with both insurers and policyholders valuing terminal wealth using exponential utility functions. However, these studies typically assume that the loss distribution is known. We extend the model proposed by Bernard et al., (2020) to account for ambiguity regarding the true loss distribution when only a limited number of samples are available.
@@ 3 Properties of entropic risk measure
The entropic risk measure is a convex-law invariant risk measure, thus satisfying the following definition.
@@
Let $\mathcal{L}^{p}:=\mathcal{L}^{p}(\Omega,\mathcal{F},{\mathbb{P}})$ denote the space of real-valued measurable functions, $X:\Omega\rightarrow\mathbb{R}$ such that ${\mathbb{E}}[|X|^{p}]<\infty$, where $p\in[1,\infty]$ and $(\Omega,\mathcal{F},{\mathbb{P}})$ is a probability space. A functional $\rho\colon\mathcal{L}^{p}\rightarrow\mathbb{R}$ is a convex law-invariant risk measure if
@@ 4. $(d)$
    $\rho(X)=\rho(X^{\prime})$ if $X\sim X^{\prime}$ for all $X,X^{\prime}\in\mathcal{L}^{p}$.
Condition $(a)$, cash-invariance property, states that $m$ is the minimum amount that should be added to a risky position to make it acceptable to a regulator. Condition $(b)$, ensures monotonicity, meaning lower losses are preferable. Condition $(c)$, convexity, ensures that diversification reduces risk. Lastly, condition $(d)$, law invariance, states that two random variables with the same distribution should have equal risk.
To ensure that the mean and variance of $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$ are finite, we make the following assumption:
@@
${\mathbb{P}}(|\ell({{\boldsymbol{\xi}}})|>a)\leq G\exp(-a\alpha C)\quad\text{ for all }a\geq 0,$
for some $G>0$ and $C>2$.
Assumption 1 further restricts the space of loss functions in Definition 1. We consider loss functions $\ell({{\boldsymbol{\xi}}})\in\mathcal{L}^{p},\,p\in[1,\infty)$ for which Assumption 1 holds. Although Assumption 1 might appear restrictive, intuitively we cannot expect to correct the bias with finite number of samples if the true risk is $\infty$. For many light tail distributions, like the GMM, Assumption 1 holds, thus entropic risk is finite for any finite risk aversion parameter $\alpha$.
@@
4 Bias correction using bias-aware bootstrapping
In this section, we introduce our proposed estimators designed to address the underestimation problem associated with the empirical entropic risk estimator, $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))$. The true bias is given by $\delta=\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))-{\mathbb{E}}_{\mathcal{D}_{N}}[\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))]$. Since the true distribution ${\mathbb{P}}$ is unknown, the exact bias cannot be determined. A typical approach in the literature to estimate the bias is using bootstrapping which samples repeatedly from the empirical distribution. Such bootstrapping procedure has been shown to be weakly consistent (DasGupta,, 2008), however, it exhibits significant bias for small sample sizes. In this paper, we propose a modification to the bootstrap algorithm, namely, we first fit a distribution $\mathbb{Q}_{N}$ using the $N$ i.i.d. loss scenarios $\mathcal{X}:=\{\ell({{\boldsymbol{\xi}}}_{1}),\ell({{\boldsymbol{\xi}}}_{2}),\cdots,\ell({{\boldsymbol{\xi}}}_{N})\}$ and then repeatedly sample from $\mathbb{Q}_{N}$, instead of resampling from the empirical distribution. We will demonstrate that fitting a distribution is crucial step in reducing bias. However, merely fitting a distribution using maximum likelihood estimation (MLE) does not fully resolve the underestimation issue, which is why we introduce bias-aware procedures to better fit the data, see Sections 4.1 and 4.2.
Let a sample $\zeta\sim\mathbb{Q}_{N}$ capture the loss $\ell({{\boldsymbol{\xi}}})$ associated with the uncertain parameter ${{\boldsymbol{\xi}}}$. Similar to Assumption 1, the following assumption ensures that the mean and variance of $\exp(\alpha\zeta)$ are finite.
@@ 4 Bias correction using bias-aware bootstrapping
Suppose that the tails of $\zeta\sim\mathbb{Q}_{N}$ are almost surely uniformly exponentially bounded. Namely, with probability one with respect to the sample and the fitting procedure of $\mathbb{Q}_{N}$, there exists some $G>0$ and some $C>2$ such that $\zeta$ satisfies Assumption 1 for all $N\geq 1$.
This assumption is not limiting since we have assumed that $\ell({{\boldsymbol{\xi}}})$ satisfies this assumption under ${\mathbb{P}}$. In practice, this assumption could be satisfied by properly defining the set of models used to fit to the observed realizations of $\ell({{\boldsymbol{\xi}}})$.
Next we introduce the necessary notation to describe our estimation procedure. Let $\rho_{\mathbb{Q}_{N}}(\zeta)$ denote the entropic risk for the distribution $\mathbb{Q}_{N}$. Let ^$\hat{\mathbb{Q}}_{N,N}$ represent the empirical distribution of $N$ values drawn i.i.d. from the estimated distribution $\mathbb{Q}_{N}$, and $\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)$ denote the corresponding empirical entropic risk. Notice that since $\mathbb{Q}_{N}$ was estimated using $N$ loss scenarios, it is itself random, thus $\rho_{\mathbb{Q}_{N}}(\zeta)$ and $\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)$ are random variables as well. The proposed estimator for the bias of the entropic risk is given by
$\delta_{N}(\mathbb{Q}_{N}):=\rho_{\mathbb{Q}_{N}}(\zeta)-\mbox{median}[\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)|\mathbb{Q}_{N}].$ (5)
Algorithm 1 provides an estimate for the bias. Given $N$ loss scenarios $\mathcal{X}$, the algorithm first estimates $\mathbb{Q}_{N}$, and then repeatedly samples $N$ i.i.d. scenarios from $\mathbb{Q}_{N}$ to form the empirical distribution ^$\hat{\mathbb{Q}}_{N,N}$. For each of the $M$ repetitions, it estimates the entropic risk, denoted by the sequence $\rho_{1},\rho_{2},\cdots,\rho_{M}$. Finally, the bias is estimated through $\hat{\delta}_{N}(\mathbb{Q}_{N})=\rho_{\mathbb{Q}_{N}}(\zeta)-\mbox{median}\{\rho_{n}\}_{n=1}^{M}$. Note that as $M$ increases, the resampling (simulation) error in the bootstrap estimate decreases, and the bootstrap estimate $\hat{\delta}_{N}(\mathbb{Q}_{N})$ converges to the true estimate $\delta_{N}(\mathbb{Q}_{N})$.
Algorithm 1 Bootstrap
1:function Bootstrap($\mathcal{X},M$)
2:     $\mathbb{Q}_{N}\leftarrow$ Fit a distribution to the loss scenarios $\mathcal{X}$
@@
7:     $\hat{\delta}_{N}(\mathbb{Q}_{N})\leftarrow\rho_{\mathbb{Q}_{N}}(\zeta)-\mbox{median}[\{\rho_{n}\}_{n=1}^{M}$]
@@
The estimator $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta_{N}(\mathbb{Q}_{N})$ is strongly asymptotically consistent.
@@
The proof involves two key steps: The first step is to establish that the empirical entropic risk converges to the true risk almost surely. This is achieved by using the strong law of large numbers and the continuous mapping theorem. The strong law of large numbers ensures that the average of exponentiated losses converges almost surely to its expected value, and the continuous mapping theorem extends this convergence to the logarithmic transformation involved in the entropic risk, leading to the almost sure convergence of the empirical risk to the true risk. The second step involves showing that the bias term, $\delta_{N}(\mathbb{Q}_{N})$ converges to zero almost surely. This is accomplished by showing that the bias, calculated under any sequence of fitted distributions {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that satisfy Assumption 2 can be made arbitrarily small for sufficiently large $N$. We show that the almost sure convergence of bias to zero is equivalent to proving that the median of a random variable $X_{N}$ representing the ratio of empirical and true risk under ¯$\bar{\mathbb{Q}}_{N}$ converges to $1$ almost surely. To establish this, Chebyshev’s inequality is used rather than the CLT, as it applies for any finite $N\geq 1$. By setting the upper tail probability to 25% in the Chebyschev’s inequality, the median of $X_{N}$ is bounded within an interval around $1$ which becomes smaller as $N$ increases. Consequently, the median of $X_{N}$ converges to $1$ almost surely. Thus, for any sequence of distributions {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ which satisfy Assumption 2, the bias converges almost surely to zero.
Our approach does not rely on asymptotic or parametric (Gaussian, for instance) assumptions made in the literature to correct the bias. Many bias correction methods assume that the estimator’s error is asymptotically normally distributed and linear, with the error term diminishing faster than one over the square root of the sample size. This allows for deriving a bias correction term (see Iyengar et al., (2023)). However, these corrections hold only asymptotically. In contrast, our method is motivated by addressing the underestimation issue in finite samples. While bias estimated via bootstrapping can be added to the empirical risk to construct a strongly asymptotically consistent estimator, a limited sample size can introduce statistical errors when using MLE to estimate the distribution, leading to inaccurate bias corrections.
@@
Among the options for choosing the distribution $\mathbb{Q}_{N}$, we utilize a $\text{GMM}(\boldsymbol{\pi},\boldsymbol{\mu},\boldsymbol{\sigma})$ where $\boldsymbol{\pi}\in\mathbb{R}^{Y}$ denotes the weights of the $Y$ mixtures, and $\boldsymbol{\mu}\in\mathbb{R}^{Y}$ and $\boldsymbol{\sigma}\in\mathbb{R}^{Y}$ denote the means and standard deviations of the mixtures, respectively. There are many advantages for using GMM: $(i)$ GMMs are universal density approximators, meaning they can approximate any smooth density given sufficient data, $(ii)$ the moment-generating function of a random variable $\zeta\sim\text{GMM}({{\boldsymbol{\pi}}},{{\boldsymbol{\mu}}},{{\boldsymbol{\sigma}}})$ exists for all $\alpha$, and thus the entropic risk $\rho_{\mathbb{Q}_{N}}(\zeta)=\frac{1}{\alpha}\log\left(\sum_{y=1}^{Y}\pi_{y}\exp(\alpha\mu_{y}+\frac{\alpha^{2}}{2}\sigma_{y}^{2})\right)$ can be obtained in closed form. This eliminates the need to estimate the entropic risk through simulation, which can be particularly beneficial when the risk aversion parameter $\alpha$ is large, as this would otherwise require a large number of samples for accuracy.
@@
The natural approach for fitting a GMM is to use MLE, typically achieved via the Expectation-Maximization algorithm (Dempster et al.,, 1977). In the following example, we demonstrate that our bootstrapping procedure, which fits a GMM using the Expectation-Maximization algorithm, still underestimates the true risk.
@@
Figure 2: Bias correction estimated from fitting a GMM by MLE followed by bootstrapping.
Bootstrapping corrects bias in empirical entropic risk estimator if the bias in $N$ i.i.d. samples from the fitted distribution matches that of the true distribution ${\mathbb{P}}$. We can satisfy this condition if we can correctly estimate the true underlying distribution ${\mathbb{P}}$. Although GMM is a universal approximator of density, MLE with limited samples introduces statistical errors in parameter estimation. This motivates the use of a “bias-aware” distribution matching algorithm (described in the next section) for learning a GMM $\mathbb{Q}_{N}$, whose $N$ i.i.d. samples capture the bias present in $N$ i.i.d samples from the true distribution ${\mathbb{P}}$. This approach is inspired by “decision-aware” learning of estimators for contextual optimization problems (Elmachtoub and Grigas,, 2021; Donti et al.,, 2017; Sadana et al.,, 2024) where one trades statistical accuracy for decision quality.

### v3

@@ 2.1 Risk estimation
Quantitative risk measurement often relies on precise estimation of risk measures commonly used in finance and actuarial science (McNeil et al.,, 2005). Calculating risk for multidimensional random variables can be challenging due to the need for complex integrals, often approximated using Monte Carlo simulation. When the underlying distribution isn’t directly accessible and only limited samples are available, Monte Carlo-based risk estimators tend to underestimate the actual risk. Kim and Hardy, (2007) addressed this by using bootstrapping to correct bias in Value at Risk (VaR) and Conditional Tail Expectation (CTE) estimates, with Kim, (2010) extending this approach to general distortion risk measures. Bootstrapping involves sampling with replacement from the empirical distribution, calculating the statistic, and averaging the outcomes over multiple iterations. However, bootstrap estimates still tend to underestimate risk in finite samples due to the lack of extreme tail scenarios (see Figure 4). In contrast, our bootstrapping procedure draws samples from a “bias-aware” fitted distribution that better accounts for tail scenarios.
Several approaches have been proposed in the literature for estimating tail risks. Lam and Mottet, (2017) introduce a distributionally robust optimization based method to construct worst-case bounds on tail risk, assuming the density function is convex beyond a certain threshold. Extreme Value Theory (EVT) is commonly used to estimate tail risk measures, such as CVaR, by fitting a Generalized Pareto Distribution to values exceeding a threshold. Troop et al., (2021) develop an asymptotically unbiased CVaR estimator by correcting the bias in the estimates obtained via maximum likelihood. In Kuiper et al., (2024), the authors derive DRO-based estimators for EVT statistics to account for model misspecification due to scenarios outside the asymptotic tails. While these methods focus on upper-tail risk, they are not directly applicable to the entropic risk measure. Instead of fitting an extreme value distribution, we utilize a parametric two-component Gaussian Mixture Model (GMM), which provides a closed-form expression for the entropic risk and captures both the mean and the tails of the data.
@@ 2.2 Correcting optimistic bias
Several methods have been proposed to correct the bias of SAA in linear optimization problems under the assumption that the true data distribution is Gaussian. For example, Ito et al., (2018) introduce a perturbation technique that generates parameters around the true values under Gaussian error assumptions to achieve an asymptotically unbiased estimator of the true loss. Similarly, Gupta et al., (2024) derive estimators for the out-of-sample performance of in-sample optimal policies under a Gaussian distribution and offer extensions to approximate Gaussian cases, leveraging the structure of linear optimization problems. However, extending these methods to our nonlinear problem is challenging. Moreover, our objective is to find optimal policies with low out-of-sample risk rather than merely estimating this risk. Since CV risk could be biased estimate of the out-of-sample risk, we employ our bias correction procedure to correct it, enabling appropriate calibration of the regularization parameter. Importantly, our approach does not rely on Gaussian assumptions for the uncertain parameter or the structure of the objective function.
In Siegel and Wagner, (2023), the authors analytically characterize the bias in SAA policies for a data-driven newsvendor problem, providing an asymptotically debiased profit estimator by leveraging the asymptotic properties of order statistics. Iyengar et al., (2023) introduce an Optimizer’s Information Criterion (OIC) to correct bias in SAA policies, generalizing the approach by Siegel and Wagner, (2023) to other loss functions. However, OIC requires access to the gradient, Hessian, and influence function of the decision rule, which can be challenging to obtain in general constrained optimization problems. Moreover, the form of optimal policy is known for risk-neutral newsvendor problems but not for entropic risk minimization problems.
@@ 2.3 Insurance pricing
The design of insurance contracts has been widely studied since the foundational work of Arrow (Arrow,, 1963, 1971). Under the assumption that premiums are proportional to the policy’s actuarial value, it has been shown that an expected utility-maximizing policyholder will choose full coverage above a deductible. Various extensions of Arrow’s model have been proposed to account for the risk aversion of both the insured and the insurer, using criteria such as mean-variance (Kaluszka, 2004a, ; Kaluszka, 2004b, ), Value at Risk (VaR), and Tail VaR (Cai et al.,, 2008). Bernard and Tian, (2010) incorporate regulatory constraints on the insurer’s insolvency risk through VaR. Cheung et al., (2014) extend these models to multiple policyholders with fully dependent risks (comonotonicity), where the insurer utilizes convex law-invariant risk measures. Bernard et al., (2020) further explore different levels of dependence among policyholders, with both insurers and policyholders using exponential utility functions. However, these studies typically assume that the loss distribution is known. We extend the model proposed by Bernard et al., (2020) to account for ambiguity regarding the true loss distribution when only a limited number of samples are available.
@@ 3 Properties of entropic risk measure
Let $(\Omega,\mathcal{F},{\mathbb{P}})$ is a probability space and let $\mathcal{L}^{p}:=\mathcal{L}^{p}(\Omega,\mathcal{F},{\mathbb{P}})$ denote the space of real-valued measurable functions, $X:\Omega\rightarrow\mathbb{R}$ such that ${\mathbb{E}}[|X|^{p}]<\infty$, for some $p\geq 1$. The entropic risk measure is a convex, law invariant risk measure (Föllmer and Schied,, 2002), thus satisfying the following definition.
@@
A functional $\rho\colon\mathcal{L}^{p}\rightarrow\bar{\mathbb{R}}$, where $\bar{\mathbb{R}}:=\mathbb{R}\cup\{\infty\}$, is a convex law-invariant risk measure if
@@ 4. $(d)$
    $\rho(X)=\rho(X^{\prime})$ for all $X,X^{\prime}\in\mathcal{L}^{p}$ such that $X=X^{\prime}$ in distribution.
Condition $(a)$, also known as cash-invariance property, states that $m$ is the minimum amount that should be added to a risky position to make it acceptable to a regulator. Condition $(b)$, ensures monotonicity, meaning lower losses are preferable. Condition $(c)$, convexity, ensures that diversification reduces risk. Lastly, condition $(d)$, law invariance, states that two random variables with the same distribution should have equal risk.
Letting ${{\boldsymbol{\xi}}}:\Omega\rightarrow\mathbb{R}^{d}$ be a random vector and $\ell({{\boldsymbol{\xi}}})\in\mathcal{L}^{p}$ the random loss that it produces, we will further impose the following assumption to ensure that the entropic risk of $\ell({{\boldsymbol{\xi}}})$ is finite, together with the mean and variance of $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$.
@@
${\mathbb{P}}(|\ell({{\boldsymbol{\xi}}})|>a)\leq G\exp(-a\alpha C),\quad\forall a\geq 0,$
for some $G>0$ and $C>2$. Equivalently, the moment-generating function ${\mathbb{E}}[\exp(t\ell({{\boldsymbol{\xi}}}))]\in\mathbb{R}$ for all $t\in[-\alpha C,\alpha C]$ for some $C>2$, see Lemma 9 in Appendix A.8 for a proof of equivalence.
Assumption 1 further restricts the space of loss functions in Definition 1 in order to work with random variables that are “well-behaved” from the point of view of entropic risk estimation at a risk tolerance level of $\alpha$. Indeed, our assumptions will ensure that the empirical estimator is asymptotically consistent. We note that our assumption relates to $\mathcal{L}_{M}$, the set of random variables with finite-valued moment generating functions, through the following inclusion: $\mathcal{L}^{\infty}\subseteq\mathcal{L}_{M}\subseteq\mathcal{L}_{\alpha}\subseteq\mathcal{L}^{p}$ with $\mathcal{L}_{\alpha}$ as the set of random variables in $\mathcal{L}_{p}$ that satisfy Assumption 1.
@@
4 Bias mitigation using bias-aware bootstrapping
In this section, we introduce our proposed estimators designed to address the underestimation problem associated with the empirical entropic risk estimator, $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))$. The true bias is given by $\delta=\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))-{\mathbb{E}}[\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))]$, where the expectation is taken with respect to the randomness in $\mathcal{D}_{N}$. Since the true distribution ${\mathbb{P}}$ is unknown, the exact bias cannot be determined. Several approaches rely on CLT to devise asymptotically unbiased estimators. However, with heavy-tailed losses (high risk aversion), a large number of samples is required before the estimator’s error tails exhibit Gaussian behavior (see Figure 1). For instance, a typical approach in the literature to devise an unbiased estimator is to use bootstrapping which samples repeatedly from the empirical distribution. Such bootstrapping procedure has been shown to be weakly consistent (DasGupta,, 2008), however, it exhibits significant bias for small sample sizes. In this paper, we propose a modification to the classical bootstrap algorithm, namely, we first fit a distribution $\mathbb{Q}_{N}$ using the $N$ i.i.d. loss scenarios $\mathcal{S}:=\{\ell({{\boldsymbol{\xi}}}_{1}),\ell({{\boldsymbol{\xi}}}_{2}),\cdots,\ell({{\boldsymbol{\xi}}}_{N})\}$ and then repeatedly sample from $\mathbb{Q}_{N}$, instead of resampling from the empirical distribution. We will demonstrate that fitting a distribution is a crucial step in reducing bias in finite samples, while still ensuring that the estimator is strongly asymptotic consistent. Nevertheless, merely fitting $\mathbb{Q}_{N}$ using MLE does not fully resolve the underestimation issue, which is why we introduce bias-aware procedures to better fit the data, see Sections 4.1 and 4.2.
Let $\zeta\sim\mathbb{Q}_{N}$ capture the loss $\ell({{\boldsymbol{\xi}}})$ associated with the uncertain parameter ${{\boldsymbol{\xi}}}$. Similar to Assumption 1, the following assumption ensures that the mean and variance of $\exp(\alpha\zeta)$ are finite.
@@ 4 Bias mitigation using bias-aware bootstrapping
Suppose that the tails of $\zeta\sim\mathbb{Q}_{N}$ are almost surely uniformly exponentially bounded. Namely, with probability one with respect to the sample and the fitting procedure of $\mathbb{Q}_{N}$, there exists some $G>0$ and some $C>2$ such that $\zeta\sim\mathbb{Q}_{N}$ satisfies Assumption 1 for all $N\geq 1$.
This assumption is not limiting since we have assumed that $\ell({{\boldsymbol{\xi}}})$ satisfies this assumption under ${\mathbb{P}}$. In practice, this assumption could be satisfied by properly defining the set of models used to estimate $\mathbb{Q}_{N}$ from the loss scenarios $\mathcal{S}$.
Next we introduce the necessary notation to describe our estimation procedure. Let $\rho_{\mathbb{Q}_{N}}(\zeta)$ denote the entropic risk for the distribution $\mathbb{Q}_{N}$. Let ^$\hat{\mathbb{Q}}_{N,N}$ represent the empirical distribution of $N$ values drawn i.i.d. from the estimated distribution $\mathbb{Q}_{N}$, and $\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)$ denote the corresponding empirical entropic risk. Notice that since $\mathbb{Q}_{N}$ was estimated using $N$ loss scenarios, it is itself random, thus $\rho_{\mathbb{Q}_{N}}(\zeta)$ and $\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)$ are random variables as well. Our proposed estimator for the bias of the entropic risk is given by
$\delta_{N}(\mathbb{Q}_{N}):=\mbox{median}\left(\rho_{\mathbb{Q}_{N}}(\zeta)-\rho_{\hat{\mathbb{Q}}_{N,N}}(\zeta)|\mathbb{Q}_{N}\right),$ (5)
where the median is taken with respect to randomness of samples from $\mathbb{Q}_{N}$. Algorithm 1 provides an estimate for the bias. Given $N$ loss scenarios $\mathcal{S}$, the algorithm first estimates $\mathbb{Q}_{N}$, and then repeatedly samples $N$ i.i.d. scenarios from $\mathbb{Q}_{N}$ to form the empirical distribution ^$\hat{\mathbb{Q}}_{N,N}$. For each of the $M$ repetitions, it estimates the entropic risk, denoted by the sequence $\rho_{1},\rho_{2},\cdots,\rho_{M}$. Finally, the bias is estimated through $\hat{\delta}_{N}(\mathbb{Q}_{N})=\mbox{median}\{\rho_{\mathbb{Q}_{N}}(\zeta)-\rho_{n}\}_{n=1}^{M}$. Note that as $M$ increases, the resampling (simulation) error in the bootstrap estimate decreases, and the bootstrap estimate $\hat{\delta}_{N}(\mathbb{Q}_{N})$ converges to the true estimate $\delta_{N}(\mathbb{Q}_{N})$.
Algorithm 1 Bootstrap bias correction
1:function BootstrapBiasCorrection($\mathcal{S},M$)
2:     $\mathbb{Q}_{N}\leftarrow$ Fit a distribution to the loss scenarios $\mathcal{S}$
@@
7:     $\hat{\delta}_{N}(\mathbb{Q}_{N})\leftarrow\mbox{median}[\{\rho_{\mathbb{Q}_{N}}(\zeta)-\rho_{n}\}_{n=1}^{M}$]
@@
Under Assumptions 1 and 2, the estimator $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta_{N}(\mathbb{Q}_{N})$ is strongly asymptotically consistent.
@@
The proof involves two key steps: The first step is to establish that the empirical entropic risk converges to the true risk almost surely. This is achieved by using the strong law of large numbers and the continuous mapping theorem. The strong law of large numbers ensures that the average of exponentiated losses converges almost surely to its expected value, and the continuous mapping theorem extends this convergence to the logarithmic transformation involved in the entropic risk, leading to the almost sure convergence of the empirical risk to the true risk. The second step involves showing that the bias term, $\delta_{N}(\mathbb{Q}_{N})$, converges to zero almost surely. This is accomplished by showing that the bias, calculated under any sequence of fitted distributions {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that satisfy Assumption 2 can be made arbitrarily small for sufficiently large $N$. We show that the almost sure convergence of bias to zero is equivalent to proving that the median of a random variable $X_{N}$ representing the ratio of empirical and true risk under ¯$\bar{\mathbb{Q}}_{N}$ converges to $1$ almost surely. To establish this, Chebyshev’s inequality is used rather than the CLT, as it applies for any finite $N\geq 1$. By setting the upper tail probability to 25% in the Chebyschev’s inequality, the median of $X_{N}$ is bounded within an interval around $1$ which becomes smaller as $N$ increases. Consequently, the median of $X_{N}$ converges to $1$ almost surely. Thus, for any sequence of distributions {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ which satisfy Assumption 2, the bias $\delta_{N}(\mathbb{Q}_{N})$ converges almost surely to zero.
Our approach does not rely on asymptotic or parametric (Gaussian, for instance) assumptions made in the literature to correct the bias. Due to the heavy tailed loss, the asymptotically unbiased estimators proposed in the literature underestimate the risk in the finite sample case. In fact, the underestimation problem persists in our proposed method as well if we naively use our bootstrapping procedure in finite samples, e.g., using MLE to fit a distribution. Since our procedure is flexible in the choice of distributions to fit to the data and the fitting procedure, we will account for the bias due to the tail scenarios in the following subsections.
@@
Among the options for choosing the distribution $\mathbb{Q}_{N}$, we utilize a Gaussian Mixture Model (GMM), $\mathbb{Q}^{{\boldsymbol{\theta}}}$, with parameters ${{\boldsymbol{\theta}}}:=(\boldsymbol{\pi},\boldsymbol{\mu},\boldsymbol{\sigma})$, where $\boldsymbol{\pi}\in\mathbb{R}^{Y}$ denotes the weights of the $Y$ mixtures, and $\boldsymbol{\mu}\in\mathbb{R}^{Y}$ and $\boldsymbol{\sigma}\in\mathbb{R}^{Y}$ denote the means and standard deviations of the mixtures, respectively. There are two advantages for using GMM. First, GMMs are universal density approximators, meaning they can approximate any smooth density given sufficient data, and second, the moment-generating function of a random variable $\zeta\sim\mathbb{Q}^{{\boldsymbol{\theta}}}$ exists for all $\alpha$, and thus the entropic risk $\rho_{\mathbb{Q}^{{\boldsymbol{\theta}}}}(\zeta)=\frac{1}{\alpha}\log\left(\sum_{y=1}^{Y}\pi_{y}\exp(\alpha\mu_{y}+\frac{\alpha^{2}}{2}\sigma_{y}^{2})\right)$ can be obtained in closed form. This eliminates the need to estimate the entropic risk through simulation in step 7 of Algorithm 1, which can be particularly beneficial when the risk aversion parameter $\alpha$ is large, as this would otherwise require a large number of samples for accuracy.
@@
The natural approach for fitting the parameters ${{\boldsymbol{\theta}}}$ of a GMM is to use MLE, typically achieved via the Expectation-Maximization (EM) algorithm (Dempster et al.,, 1977). The following example demonstrates that with limited samples, if step 2 of Algorithm 1 utilizes the Expectation-Maximization algorithm to fit a GMM, the estimate will still underestimate the true risk.
@@
Figure 2: Statistics of bias correction estimated from fitting a GMM by MLE followed by bootstrapping over $100$ resampling from the underlying distribution.
We next propose different distribution fitting strategies that can aid the bias mitigation in finite samples. The aim is to fit a distribution such that $N$ i.i.d. samples drawn from the fitted distribution replicate the bias observed in $N$ i.i.d. samples from the true distribution ${\mathbb{P}}$. We refer to this approach as “bias-aware” distribution matching. This concept is inspired by “decision-aware” learning methods in contextual optimization problems (Elmachtoub and Grigas,, 2021; Donti et al.,, 2017; Grigas et al.,, 2023; Sadana et al.,, 2025), where statistical accuracy is deliberately traded for improved decision outcomes.